OpenStudy (anonymous):

solve to get x : log⁡(x^2+1)=1

OpenStudy (anonymous):

the only way for \(\log(x^2+1)=1\) is if \(x^2+1=10\)

OpenStudy (anonymous):

assuming of course that than means log base ten \[\log_{10}(A)=B\iff A=10^B\]

OpenStudy (anonymous):

A little confused... would it be x^2+1=log10(1) ? ....

OpenStudy (anonymous):

noooo \[\Large \log(x^2+1)=1\\\Large 10^{\log(x^2+1)}=10^1\\\Large x^2 +1 =10\]

OpenStudy (anonymous):

Ohhh and then x^2=10-1 -> x^2=9 and the x=sqrt(9) which then would be 3, @PeterPan ?

OpenStudy (anonymous):

3 or something else ;) And this is all assuming that the log is log base 10 ok? ^.^

OpenStudy (anonymous):

Aaaalright.... -.- I'll figure it out hopefully. Thanks anyways, @PeterPan

OpenStudy (anonymous):

besides, the more proper way of solving x^2 - 9 = 0 is by factoring ^.^ That gives you the two solutions you need, yes?

OpenStudy (anonymous):

I need two solutions? The question just says, solve the functions:/

OpenStudy (anonymous):

Yes, but there are two possible answers to x^2 - 9 = 0 Did you try -3? I didn't think so :P Silly Sulle :D

OpenStudy (anonymous):

Ahh it could be. But hell. I just want to get it as close to the right answer as possible. this'll do :P Thanks though, @PeterPan ! :D Love the username

OpenStudy (anonymous):

It'd be better if we could be a little more sure, though... Oh well, guess it can't be helped~ see you around :3